OpenStudy (anonymous):

Gunther drew a circle. The radius of the circle is 20 inches. He uses the formula below to determine the area of his circle. A = Pi r 2 What is the area, in square inches, of gunthers circle? Leave your answer in terms of Pi Choices: A. 10 pi B. 40 pi C. 100 D. 400 pi Explain how you determined your answer.

OpenStudy (anonymous):

D.

OpenStudy (anonymous):

A=pir^2. A=pi(20)^2/ A=400pi (inches)

OpenStudy (anonymous):

Got it?